OpenStudy (anonymous):

Look at the formula for the area of a circle below. A = pi r2

OpenStudy (anonymous):

\[A=\pi r ^{2}\]

OpenStudy (anonymous):

\[ r = \sqrt{\frac{A}{\pi}} \]

OpenStudy (anonymous):

Which of the following equations can be used to solve for the radius (r)?

OpenStudy (anonymous):

See above.